11.1.40 A physical therapist wants to determine the difference in the proportion of men and women who participate in regular sustained physical activity. What sample size should be obtained if she wishes the estimate to be within five percentage points with 99​% ​confidence, assuming that ​(a) she uses the estimates of 22.6​% male and 19.2​% female from a previous​ year? ​(b) she does not use any prior​ estimates?
Answer #1

solution:-

given that error E = 5% = 0.05

the value of 99% confidence from z table is 2.58

(a) here given that p1 = 0.226 , p2 = 0.192

formula for sample size

=> n = (p1(1-p1) + p2(1-p2))*(z/E)^2

=> n = (0.226(1-0.226)+0.192(1-0.192))*(2.58/0.05)^2

=> n = 879


​(b) she does not use any prior​ estimates

we have to take p1 = 0.5 , p2 = 0.5

=> n = (0.5(1-0.5)+0.5(1-0.5))*(2.58/0.05)^2

=> n = 1332